OpenStudy (anonymous):

*Logarithmic Function* A certain radioactive substance decays exponentially.The percent , P , of the substance left after , t, years is given by formula P(t)=100(1.04)^-t.Determine the instantaneous rate of decay of the substance when it has reached its half-life .

OpenStudy (anonymous):

something wrong here \[P(t)=100(1.04)^{-1}\] doesn't have a \(t\) in it

OpenStudy (anonymous):

I'm sorry, i meant the power, -1 as the variable -t, thank you.

OpenStudy (anonymous):

\[P(t)=100(1.04)^{-t}\] like this?

OpenStudy (anonymous):

Yes, but I was able to solve it now, I had a mindlapse, sorry, you can confirm is the answer I got, t = 17.67 years is correct?

OpenStudy (anonymous):

well you need two things, first the time when it is half, and then the instantaneous rate of change

OpenStudy (anonymous):

first one you solve using logs \[\frac{1}{2}=(1.04)^{-t}\] \[t=-\frac{\ln(.5)}{\ln(1.04)}\]

OpenStudy (anonymous):

which is t = 17.67, then find the derivative of the given equation?

OpenStudy (anonymous):

yeah that looks right, 17.67

OpenStudy (anonymous):

yeaaa, never mind, the question I had typed had a mistake

OpenStudy (anonymous):

yes, and then replace \(t\) byu 17.67

OpenStudy (anonymous):

But thanks so much, I'm going to close the question, thanks a lot ! fast answer and excellent help.

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

after you get 17.67 how do u find it

OpenStudy (anonymous):

i got I.R.O.C -0.4035

OpenStudy (anonymous):

IS THAT THE ANSWER